OpenStudy (precal):

solve for x x=x^(1/2) +6

OpenStudy (anonymous):

7/2=0

OpenStudy (precal):

? no solution or did you mean to say x=7/2

OpenStudy (anonymous):

no solution for x ...7/2=

OpenStudy (anonymous):

show??

OpenStudy (precal):

show what? I don't understand your response

OpenStudy (precal):

I am being lazy, I really should go read that part of my book

OpenStudy (anonymous):